What is the slope intercept form of 3y+x-9=0?
Answer by chicosgurrl(2)

The formula for slope-intercept is :
.
y=mx+b
.
so, they gave you 3y+x-9=0
so make it look like the slope intercept form by getting Y on one side of the equal sign.
.
In order to do that you need to subtract x and add 9 to both sides:
3y+x-9=0
-x+9=-x+9
_________
3y=-x+9 ...but the Y must be by itself so divide both side by 3:
.
y= -1/3x+3...(final answer)
.
so that is how to put it into slope intercept form.